# Infrared-Radio-follow-up Observations for Detection of the Magnetic Radio Emission of Extra-Solar Planets: A New Window to Detect Exoplanets

### v1

@@ 2.5 Observation Strategy
2.6 Detectablity of the Planetary Radio Signal
@@
where, $I(\nu,T_{p})$ denotes the emitted power per unit area of the emitting surface, per unit solid angle, and unit frequency. Integrating it across the energy range of the W149 filter and over a hemisphere yields the thermal flux of the planet as observed by Roman. The flux of the planet resulting from the reflection ($F_{ref}$) of the parent star’s light is determined by
@@
where $g(\Phi)$ denotes the fraction of the illuminated portion of the planet visible to the observer, $A_{g}$ represents the geometric albedo ($A_{g}=2/3~{}A_{B}$) (Bagheri et al., 2019). Consequently, the overall flux received from the planet ($F_{p}$) can be calculated as the combined sum of the thermal flux and the reflection flux. In the final stage, we add realistic noise. The total noise consists of the intrinsic Poisson fluctuation in the flux received from the microlensing event in a single exposure, the intrinsic Poisson fluctuation within the point spread function (PSF) representing the background sky flux in a single exposure, the read-out and dark noise (Bagheri et al., 2019). Finally, we determine the range within which the noisy signal exhibits magnitudes below the zero-point magnitude of Roman and above the saturation magnitude of 14.8 (Penny et al., 2019).
@@
We use the SKA image sensitivity for different ranges of frequencies as summarized in LABEL:tab:SKA. Therefore, any event with flux density greater than the SKA sensitivity for at least 4 hours of imaging is considered a detected event.
@@ 3 Results
We simulate 150,000 single-lens and binary-lens microlensing events. Our Monte Carlo simulations indicate the likelihood of detecting radio emission of rocky planets using either single-lens or binary-lens configurations is exceedingly low with the SKA telescope. Consequently, Jovian planets exhibit the highest detection probabilities in both scenarios. This outcome primarily stems from their stronger magnetic field and higher frequency of radio emission relative to rocky planets, resulting in a substantial portion of their radiation falling within the SKA sensitivity. Simulations of Roman-SKA data predict, under the assumption of single-lens configuration, that the number of detecting a Jovian planet with orbital periods of 2 days to 10 years is about 2.018 %. For the binary-lens simulation, this number increases to 9.01 % due to the high magnification of caustic crossings.
@@
Figure 3 shows our simulation’s distributions of planetary parameters. Exoplanets with detected radio fluxes are in orange. This decreasing abundance of the simulated high-mass planets with periods $<1500$ days (corresponds to $a<2.5$ au), as seen in Figure 3-a, results from the relationship between giant planets’ abundance and the host stars’ metallicity as outlined by (Johnson et al., 2010). This relationship was established for planets with $a<2.5$ au. These criteria typically lead to conservative estimates of the overall count of detected giant planets with smaller orbital periods. The planetary radio flux in terms of magnification of the microlensing events is shown in Figure 3-b. Exoplanets within the shaded region exhibit emission frequencies below 50 MHz, rendering their signals undetectable despite their radio emissions being of high intensity and possessing a significant magnification factor during microlensing events. Figures 3-d and 3-e represent the distance of the star-planet system relative to the exoplanet mass and the magnification of the microlensing events. Notably, the range extends up to 10 kpc for exoplanets with detectable signals, underscoring the importance of this observation method. These exoplanets are beyond the reach of current technologies for detection in any other wavelength bands using conventional methods. Identifying such star-planet systems can enhance our understanding of planet formation rates within the galactic bulge and facilitate the detection of extragalactic exoplanets.
@@
The detection efficiency in terms of mass and semi-major axis of the exoplanets is illustrated in Figure 4. Detected planets exhibit masses ranging from $M_{p}=0.6$ to $13.5M_{J}$ and semi-major axes spanning from $a=0.01$ to $6.9$ au.
@@
We conducted simulations targeting coordinates $l=0.5^{\circ}$ and $b=-1.5^{\circ}$. In this direction, as reported by (Penny et al., 2019), the estimated number of source stars detectable by the Roman telescope with magnitudes below 25 is $240\times 10^{6}$. The optical depth, averaged over the distances of the source stars, is $\tau=2.4\times 10^{-6}$ (Mróz et al., 2019), while our simulation yields an average value of $<\epsilon/t_{E}>=0.035$. Consequently, approximately 1155 exoplanets could be detected with Roman-SKA observations toward the Galactic bulge (Bagheri et al., 2019). However, it may not be feasible to follow up on all Roman’s detected microlensing events; nonetheless, a 10-20% follow-up of these events could significantly impact exoplanet sciences and the measurement of exoplanetary magnetic fields. These findings are summarized in Table 3.
@@
Figure 3: Characteristics of the exoplanets within the simulated event. Exoplanets exhibiting detectable radio signals are highlighted in orange.
@@ 4 Discussion
Our comprehension of planetary magnetic fields remains largely confined to the solar system. Expanding our knowledge and observation of exoplanetary magnetic fields could provide valuable insights into their internal compositions and dynamics, which are otherwise challenging to ascertain. In this study, we investigate an alternative approach to detecting exoplanets through their radio emissions magnified by microlensing events. Our findings reveal that the planetary radio emissions of approximately 5% of microlensing events could be detected through follow-up observations with Roman-SKA (equivalent to $\sim 9\%$ for binary-lens and $\sim 2\%$ for single-lens events). This finding gains significance when considering that, to date, only 210 exoplanets have been detected via microlensing events [^1] NASA Exoplanet Archive. These detections primarily occurred due to caustic crossings in binary-lens events, making follow-up observations of these planets nearly impracticable. In contrast, we consider exoplanets as companions within the source system of microlensing events, enabling us to conduct follow-up observations on the star-planet system.

### v2

@@ 2.5 Observation Strategy
2.6 Detectability of the Planetary Radio Signal
@@
where, $I(\nu,T_{p})$ denotes the emitted power per unit area of the emitting surface, per unit solid angle, and per unit frequency. Integrating it across the energy range of the W149 filter and over a hemisphere yields the thermal flux of the planet as observed by Roman. The flux of the planet resulting from the reflection ($F_{ref}$) of the parent star’s light is determined by
@@
where $g(\Phi)$ denotes the fraction of the illuminated portion of the planet visible to the observer, and $A_{g}$ represents the geometric albedo. We utilize a geometric albedo formula: $A_{g}=2/3~{}A_{B}$ (Bagheri et al., 2019). While this relationship may not hold true for all phase angles, its alteration does not significantly impact results. This is because the thermal radiation emitted by planets is much weaker than the flux of host stars in the Roman W149 filter. Thus, any changes in this relationship would only affect the total flux received by the Roman telescope at the order of $410^{-}$ 4 at most. Consequently, the overall flux received from the planet ($F_{p}$) can be calculated as the combined sum of the thermal flux and the reflection flux. In the final stage, we add realistic noise. The total noise consists of the intrinsic Poisson fluctuation in the flux received from the microlensing event in a single exposure, the intrinsic Poisson fluctuation within the point spread function (PSF) representing the background sky flux in a single exposure, the read-out and dark noise (Bagheri et al., 2019). Finally, we determine the range within which the noisy signal exhibits magnitudes below the zero-point magnitude of Roman and above the saturation magnitude of 14.8 (Penny et al., 2019).
@@
We use the SKA image sensitivity for different ranges of frequencies as summarized in Table LABEL:tab:SKA. Therefore, any event with flux density greater than the SKA sensitivity for at least 4 hours of imaging is considered a detected event.
@@ 3 Results
We simulate 150,000 single-lens and binary-lens microlensing events. Our Monte Carlo simulations indicate the likelihood of detecting radio emission of rocky planets using either single-lens or binary-lens configurations is exceedingly low with the SKA telescope. Consequently, Jovian planets exhibit the highest detection probabilities in both scenarios. This outcome primarily stems from their stronger magnetic field and higher frequency of radio emission relative to rocky planets, resulting in a substantial portion of their radiation falling within the SKA sensitivity. Simulations of Roman-SKA data predict, under the assumption of single-lens configuration, that the number of detecting a Jovian planet with orbital periods of 2 days to 10 years is about 3.54 %. For the binary-lens simulation, this number increases to 8.25 % due to the high magnification of caustic crossings.
@@
Figure 2 shows our simulation’s distributions of planetary parameters. Exoplanets with detected radio fluxes are in orange. As depicted in Figure 2-a, it’s notable that the majority of exoplanets exhibiting detectable signals are situated at orbital distances lower than 2 au. However, for planets with higher masses, the orbital distance can extend up to 8 au. The planetary radio flux in terms of magnification of the microlensing events is shown in Figure 2-b. Exoplanets within the shaded region exhibit emission frequencies below 50 MHz, rendering their signals undetectable despite their radio emissions being of high intensity and possessing a significant magnification factor during microlensing events. Figures 2-d and 2-e represent the distance of the star-planet system relative to the exoplanet mass and the magnification of the microlensing events. Notably, the range extends up to 10 kpc for exoplanets with detectable signals, underscoring the importance of this observation method. These exoplanets are beyond the reach of current technologies for detection in any other wavelength bands using conventional methods. Identifying such star-planet systems can enhance our understanding of planet formation rates within the galactic bulge and facilitate the detection of extragalactic exoplanets.
@@
The detection efficiency in terms of mass and semi-major axis of the exoplanets is illustrated in Figure 3. Detected planets exhibit masses ranging from $M_{p}=1$ to $10M_{J}$ and semi-major axes spanning from $a=0.01$ to $10$ au.
We conducted simulations targeting coordinates $l=0.5^{\circ}$ and $b=-1.5^{\circ}$. In this direction, as reported by (Penny et al., 2019), the estimated number of source stars detectable by the Roman telescope with magnitudes below 25 is $240\times 10^{6}$. Hence, the number of microlensing events that could be detected via Roman observation can be given by:
$N=\frac{2}{\pi}T_{obs}\times N_{\star}\times<\epsilon(t_{E})>\tau$ (23)
@@
where $T_{obs}$ is the observational time, $N_{\star}$ is the number of background stars detectable by Roman, $\epsilon(t_{E})$ is the Roman efficiency for detecting microlensing events with the duration of $t_{E}$, and $\tau$ is the optical depth (Bagheri et al., 2019). The optical depth, averaged over the distances of the source stars, is $\tau=2.4\times 10^{-6}$ (Mróz et al., 2019), and our simulation yields an average value of $<\epsilon/t_{E}>=0.035$. Therefore, around 27000 microlensing events can be detected by Roman observation (Penny et al., 2019). Consequently, approximately 1317 exoplanets could be detected with Roman-SKA observations toward the Galactic bulge. However, it may not be feasible to follow up on all Roman’s detected microlensing events; nonetheless, a 10-20% follow-up of these events could significantly impact exoplanet sciences and the measurement of exoplanetary magnetic fields. These findings are summarized in Table 3.
@@
Figure 2: Characteristics of the exoplanets within the simulated event. Exoplanets exhibiting detectable radio signals are highlighted in orange; a) distribution of semi-major axis as a function of planetary mass, b) planetary radio flux vs. magnification of the microlensing events, c) planetary radio flux as a function of the frequency of the planetary radio emissions, d) distance of the star-planet system vs. planetary mass, and e) distance of the star-planet system vs. magnification of the microlensing events.
@@ 4 Discussion
Our comprehension of planetary magnetic fields remains largely confined to the solar system. Expanding our knowledge and observation of exoplanetary magnetic fields could provide valuable insights into their internal compositions and dynamics, which are otherwise challenging to ascertain. The detection of radio emissions from confirmed exoplanets has been investigated in multiple studies (e.g., Bagheri et al. (2024a); Li et al. (2023). In this study, we investigate an alternative approach to detecting exoplanets through their radio emissions magnified by microlensing events. Our findings reveal that the planetary radio emissions of approximately 5% of microlensing events could be detected through follow-up observations with Roman-SKA (equivalent to $\sim 8\%$ for binary-lens and $\sim 3.5\%$ for single-lens events). This finding gains significance when considering that, to date, only 210 exoplanets have been detected via microlensing events [^1] NASA Exoplanet Archive. These detections primarily occurred due to caustic crossings in binary-lens events, making follow-up observations of these planets nearly impracticable. In contrast, we consider exoplanets as companions within the source system of microlensing events, enabling us to conduct follow-up observations on the star-planet system.